2.) Scores on the SAT form a normal distribution with a mean score of 500 and a standard deviation of 100.




a. What is the minimum score necessary to be in the top 15% of the SAT distribution?




b. Find the range of scores that defines the middle 80% of the distribution of SAT scores.




3. )The government would like to conduct a subsidy program for the lowest 5 percent of the families in terms of income. The government gathered data about family income and it’s found to be normally distributed with a mean of Php 130 000 and a standard deviation of Php 50 000. What is the cutoff income for the government program?


1
Expert's answer
2022-05-13T13:47:24-0400

2.)

a)


"P(X\le x)=1-0.15"

"P(Z\le \dfrac{x-500}{100})=0.85"

"\dfrac{x-500}{100}=1.036433"

"x=603.64"

Minimum score is 603.64

b)


"P(X<x_1)=0.1"

"P(Z< \dfrac{x_1-500}{100})=0.1"

"\dfrac{x_1-500}{100}=-1.28155"

"x_1=371.845"

"\dfrac{x_2-500}{100}=1.28155"

"x_2=628.155"

"P(371.845\le X\le 628.155)=0.8"

3.)



"P(X<x)=0.05"

"P(Z< \dfrac{x-130000}{50000})=0.05"

"\dfrac{x-130000}{50000}=-1.6449"

"x=47755"